Hello, please help me answer this problem: What quantity of 60 percent acid solution must be mixed with a 30 percent solution to produce 630 mL of a 50 percent solution? Im not too sure where to begin maybe make a table?
Answer by rothauserc(4718)

The problem states
What quantity of 60 percent acid solution must be mixed with a 30 percent solution to produce 630 mL of a 50 percent solution?
let x = amount of .60 solution
then 630 - x = .30 solution
and .50*630 = 315 ml of alcohol
then
.6x + .3(630 - x) = 315
.3x = 315 - 189
.3x = 126 and
x = 420
420ml of 60% solution and
210ml pf 30% solution